# Adaptive Evolutionary Framework for Safe, Efficient, and Cooperative Autonomous Vehicle Interactions

### v1

@@ Abstract
Index Terms:
Autonomous vehicle, evolutionary game theory, causal evaluation module, interactive driving, collision avoidance, adaptive strategy.
@@ I INTRODUCTION
As modern transportation systems present a unique set of challenges and risks, significantly impacting road safety. In Great Britain alone, there were over 1,700 fatalities and more than 29,000 serious injuries reported in 2022 [1]. Globally, road traffic crashes claim approximately 1.19 million lives each year [2]. These records reflects the critical need for improved safety measures in on-road driving. With the fast growth in autonomous vehicles (AVs), projected to be over 50 million by 2024 [3], modern traffic designs are evolving to better accommodate AVs. These updates address critical challenges in ensuring safe control and interaction between AVs [4, 5]. Furthermore, AVs have the potential to significantly reduce safety concerns caused by human errors such as distraction, and delayed reactions [6]. Additionally, AVs compute optimal decision-making solutions faster than human drivers, enhancing reaction to emergency situations [7]. In recent years, the AVs experience fast development and promote the driving safety and traffic efficiency [8, 9, 10, 11, 12, 13, 14].
AVs include three core stages for stable and safe driving, including the perception [15, 16, 17, 18, 19], decision-to-planning [20, 21, 22, 23], and control [24, 25, 26, 27, 10, 28]. It is challenging to achieve an effective driving between AVs without a centralized guiding system [29]. Without a centralized guiding system, the interaction between AVs can be complex due to the demand of maximizing their own profits. Additionally, AVs must be capable of balancing multiple demands of passengers, such as safety, efficiency [30, 31]. Some decision-making frameworks optimize traffic flow and improve safety [32, 33, 34, 35, 36]. Besides, the AVs is expected to be unselfish, which means they need to consider befitting the general traffic. Therefore, interactive driving between AVs requires balanced decision-making capabilities to avoid collisions, maintain high driving efficiency, and positive effects to publication. Some approaches are used for AVs’ interaction, such as rule-based methods, optimization-based methods, and game-based methods.
@@
Figure 2: The casual evaluation-based evolutionary game theory for safe, efficienct, and cooperative interactive driving.
Game-based approaches are well-suited for decision-making in interactive driving scenarios [44, 45, 46]. Two widely used game-based models are Nash and Stackelberg games. Stackelberg games are hierarchical models often used in scenarios involving a leader-follower dynamic. This approach has applications in autonomous vehicle platooning, where a leader vehicle dictates strategies that follower vehicles adopt. For example, interactions between lane-changing vehicles and their adjacent counterparts have been modeled using a Stackelberg game framework [47]. Such models enable adjustment of driving behaviors by varying emphasis on safety, efficiency, and comfort [32, 48]. However, Stackelberg games heavily rely on accurate predictions of surrounding vehicles driving styles, leading to significant errors when predictions are inaccurate. Additionally, Stackelberg games lack adaptability and flexibility in real driving, as real-world driving styles can deviate considerably from predefined styles.
@@ IV Simulation Results
Figure 13: The reward curves of CEGT, Nash and Stakelberg across varied initial state settings over 10 runs for Case 1.
Figure 14: The speed curves for each AV using CEGT, Nash, and Stakelberg across varied initial state settings over 10 runs for Case 1.
@@ IV-A Simulation Case 1
Fig. 9 compares the average number of collisions for four algorithms: CEGT, EGT, Nash, and Stakelberg. CEGT demonstrates a significant advantage with an average collision count of approximately $1.5$, the lowest among all algorithms, compared to EGT ($2.8$), Nash ($4.5$), and Stakelberg ($3.6$). The data points for CEGT are tightly concentrated around the mean, reflecting consistent and reliable results, whereas Nash and Stakelberg show wider distributions with extreme outliers, reaching collision counts as high as $16$ and $18$, respectively. This stark contrast highlights the inefficiency of the alternative methods in ensuring safety. CEGT’s ability to minimize collisions and maintain stability across scenarios underscores its effectiveness in cooperative and adaptive decision-making.
@@
Fig. 10 illustrates the average TTC for CEGT, Nash, and Stakelberg over time, highlighting how well each algorithm balances safety and efficiency. CEGT maintains TTC values consistently above the safety threshold, ensuring a low risk of collisions throughout the simulation. In contrast, Nash and Stakelberg frequently fall below the safety threshold, particularly around $t=2~\text{s}$, where their TTC values drop as low as $300~\text{s}$, indicating a high risk of collisions. CEGT also avoids prolonged periods of negative TTC, which occur when the rear AV’s speed is slower than the lead AV’s, resulting in inefficiencies. Its TTC values remain mostly positive, demonstrating that it maintains efficient driving speeds. In comparison, Nash and Stakelberg often exhibit negative TTC values or positive values below the safety threshold for extended periods. For instance, at $t=5~\text{s}$, Nash has an average TTC value below $100~\text{s}$, indicating poor speed control. Additionally, CEGT demonstrates stable and consistent performance, with its TTC values generally staying above the safety threshold without significant fluctuations. In contrast, Nash and Stakelberg show far less stability, with frequent and severe drops in TTC, making their performance both unsafe and inefficient.
@@
Fig. 11 illustrates the average TTC between AV 1 and AV 2 for the CEGT, Nash, and Stakelberg algorithms. CEGT demonstrates superior performance by maintaining TTC values consistently close to or above the safety threshold for most of the simulation. CEGT maintains TTC values above safety threshold on average, ensuring low collision risk. In contrast, Nash and Stakelberg frequently exhibit TTC values below the safety threshold, with Nash dropping to as low as $80~\text{s}$, particularly around $t=2~\text{s}$, indicating high collision risks and inefficiency. CEGT also shows stable behavior, with fewer fluctuations compared to the other algorithms. This stability highlights its ability to manage interactions between AV 1 and AV 2 effectively. On the other hand, Nash and Stakelberg experience significant volatility, with frequent drops in TTC below $0~\text{s}$, reflecting poor speed control and increased risk of unsafe situations.
@@
Fig. 12 illustrates the average TTC between AV 3 and AV 4 for the CEGT, Nash, and Stakelberg algorithms. CEGT demonstrates superior performance by maintaining TTC values consistently above the safety threshold for most of the simulation, ensuring a low risk of collisions. In contrast, Nash and Stakelberg frequently fall below the safety threshold for extended periods, indicating higher collision risks. CEGT also avoids prolonged periods of negative TTC values, which occur when AV 3’s speed is slower than AV 4’s, leading to inefficiencies. Moreover, CEGT’s TTC values remain mostly positive, highlighting effective and efficient speed management. In comparison, Nash and Stakelberg exhibit significant volatility with frequent negative TTC values, reflecting poor coordination and inefficiencies.
@@
Fig. 13 The illustrates the total cumulative reward over time for the CEGT, Nash, and Stakelberg algorithms. CEGT clearly outperforms the other algorithms by achieving the highest cumulative reward throughout the simulation. At $t=10~\text{s}$, CEGT reaches a reward of approximately $162,388$, significantly higher than Nash around $49,545$ and Stakelberg about 176,039-176,039. This demonstrates CEGT’s superior ability to balance safety, efficiency, and cooperativeness. CEGT also shows a steeper growth rate in cumulative reward, reflecting its effective decision-making and efficient coordination among vehicles. In contrast, Nash has a much slower growth, while Stakelberg consistently exhibits a decline in cumulative reward, reflecting inefficient performance and frequent unsafe behaviors. Additionally, CEGT maintains a smaller variability range compared to Nash and Stakelberg, as shown by the narrower shaded region. This indicates more stable performance across different initial settings. In comparison, Nash and Stakelberg exhibit larger variability, highlighting their instability and less reliable outcomes.
@@
Fig. 14 illustrates the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stakelberg algorithms across 10 runs. CEGT demonstrates clear advantages in maintaining higher and more stable speeds compared to Nash and Stakelberg. For AV 1, CEGT has higher average speed than Nash and Stakelberg at most time points after $2~\text{s}$. For AV 2 to AV 4, CEGT achieves an average speed of approximately $10~\text{m/s}$ to $15~\text{m/s}$ throughout the simulation, indicating efficient driving behavior. In contrast, Nash and Stakelberg show significantly lower average speeds, with Stakelberg often around $5~\text{m/s}$ to $10~\text{m/s}$, and Nash below $5~\text{m/s}$, highlighting inefficiencies.
@@ IV-B Simulation Case 2
Fig. 17 demonstrates the average number of collisions for four algorithms: CEGT, EGT, Nash, and Stakelberg. CEGT demonstrates a significant advantage with an average collision count of approximately $0.4$, the lowest among all algorithms, compared to EGT around $2.5$, Nash around $3$, and Stakelberg around $0.7$. The data points for CEGT are tightly concentrated around the mean, reflecting consistent and reliable results, whereas Nash and Stakelberg show wider distributions with extreme outliers between $5$ to $10$. This stark contrast highlights the inefficiency of the alternative methods in ensuring safety. CEGT’s ability to minimize collisions and maintain stability across scenarios underscores its effectiveness in cooperative and adaptive decision-making.
@@
Fig. 18 illustrates the cumulative reward profiles over time for four AVs using the CEGT, Nash, and Stakelberg algorithms. For AV 1, CEGT shows a steady growth in cumulative reward, reaching approximately $0.7\times 10^{5}$ at $t=10~\text{s}$. This lower reward compared to other AVs is due to AV 1 prioritizing maintaining a similar distance to other AVs, thereby sacrificing its self-reward for the overall system’s benefit. For AV 2, the CEGT reward curve is similar to Stakelberg; however, CEGT exhibits narrower variation, suggesting greater stability and adaptability to different initial settings of AVs. Similar trends are observed for AVs 3 and 4, where CEGT achieves higher rewards than the other algorithms. Although Stakelberg also presents a consistent upward trend, its wide variability indicates poor capability to adapt to random initial settings of AVs and maintain reliable performance.
@@
Figure 18: The reward curves of CEGT, Nash and Stakelberg across varied initial state settings over 10 runs for Case 2.
@@
Figure 19: The speed curves for each AV using CEGT, Nash, and Stakelberg across varied initial state settings, with a collision penalty of 100-100, over 10 runs for Case 2.
@@
Figure 20: The speed curves for each AV using CEGT, Nash, and Stakelberg across varied initial state settings, with a collision penalty of 200-200, over 10 runs for Case 2.
@@
Figure 21: The speed curves for each AV using CEGT, Nash, and Stakelberg across varied initial state settings, with a collision penalty of 300-300, over 10 runs for Case 2.
@@
To verify the generalization of CEGT, different algorithm settings are also tested. Fig. 19 shows the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stakelberg algorithms across 10 runs, with a collision penalty of -100. CEGT demonstrates clear advantages in maintaining higher and more stable speeds compared to Nash and Stakelberg. For AV 1, CEGT achieves a lower average speed than Nash and Stakelberg at most time points due to its cooperation-oriented decision-making. This sacrifice by AV 1 allows other AVs to achieve significantly higher speeds. For AVs 2 to 4, CEGT maintains an average speed of approximately $10~\text{m/s}$ to $15~\text{m/s}$ throughout the simulation, reflecting efficient driving behavior. In contrast, Nash and Stakelberg show significantly lower average speeds, with Stakelberg remaining below $10~\text{m/s}$ and Nash often below $5~\text{m/s}$, highlighting inefficiencies.
@@
Fig. 20 shows the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stakelberg algorithms across 10 runs, with a collision penalty of -200. For AV 1, CEGT displays a lower average speed compared to Nash and Stakelberg at most time points, as it prioritizes cooperation over individual performance. This trade-off enables other AVs to achieve substantially higher speeds. For AVs 2 to 4, CEGT consistently maintains an average speed ranging from $10~\text{m/s}$ to $15~\text{m/s}$ throughout the simulation, showcasing its ability to ensure efficient driving behavior. In comparison, Nash and Stakelberg exhibit significantly reduced average speeds, with Stakelberg staying below $10~\text{m/s}$ and Nash frequently falling under $5~\text{m/s}$, underscoring their lack of efficiency.
@@
Fig. 21 shows the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stakelberg algorithms across 10 runs, with a collision penalty of -300. The AV 1’s speed patterns show distinct characteristics across the three strategies. The CEGT approach maintains speeds between $8\text{-}12~\text{m/s}$ with moderate fluctuations, while demonstrating higher stability compared to other methods. The Nash equilibrium exhibits a sharp decline in speed after the initial period, settling at considerably low speeds around $2\text{-}3~\text{m/s}$. The Stackelberg approach achieves intermediate performance, maintaining speeds around $10\text{-}12~\text{m/s}$ with relatively consistent behavior. For AVs 2 through 4, similar patterns emerge but with notable differences. CEGT consistently demonstrates superior performance, maintaining speeds between $10\text{-}15~\text{m/s}$ across all vehicles, though with visible oscillations in the velocity profiles. The Nash equilibrium consistently performs poorly across all AVs, with speeds dropping dramatically to near-zero levels after the initial phase. The Stackelberg strategy shows intermediate performance but exhibits greater variance as indicated by the wider shaded regions. The shaded areas indicate that CEGT maintains relatively consistent behavior across multiple simulations, while Nash and Stackelberg strategies show higher variability. This suggests that CEGT not only achieves better average performance but also provides more reliable and predictable vehicle behavior.

### v2

@@ I INTRODUCTION
As modern transportation systems present a unique set of challenges and risks, significantly impacting road safety. In Great Britain alone, there were over 1,700 fatalities and more than 29,000 serious injuries reported in 2022 [8]. Globally, road traffic crashes claim approximately 1.19 million lives each year [12]. These records reflects the critical need for improved safety measures in on-road driving. With the fast growth in autonomous vehicles (AVs), projected to be over 50 million by 2024 [18], modern traffic designs are evolving to better accommodate AVs. These updates address critical challenges in ensuring safe control and interaction between AVs [7, 11]. Furthermore, AVs have the potential to significantly reduce safety concerns caused by human errors such as distraction, and delayed reactions [2]. Additionally, AVs compute optimal decision-making solutions faster than human drivers, enhancing reaction to emergency situations [27]. In recent years, the AVs experience fast development and promote the driving safety and traffic efficiency [26, 24, 22].
AVs include three core stages for stable and safe driving, including the perception, decision-to-planning, and control. It is challenging to achieve an effective driving between AVs without a centralized guiding system. Without a centralized guiding system, the interaction between AVs can be complex due to the demand of maximizing their own profits. Additionally, AVs must be capable of balancing multiple demands of passengers, such as safety, efficiency [25, 4]. Some decision-making frameworks optimize traffic flow and improve safety [15, 23, 3, 32, 30]. Besides, the AVs is expected to be unselfish, which means they need to consider befitting the general traffic. Therefore, interactive driving between AVs requires balanced decision-making capabilities to avoid collisions, maintain high driving efficiency, and positive effects to publication. Some approaches are used for AVs’ interaction, such as rule-based methods, optimization-based methods, and game-based methods.
@@
Game-based approaches are well-suited for decision-making in interactive driving scenarios. Two widely used game-based models are Nash and Stackelberg games. Stackelberg games are hierarchical models often used in scenarios involving a leader-follower dynamic. This approach has applications in autonomous vehicle platooning, where a leader vehicle dictates strategies that follower vehicles adopt. For example, interactions between lane-changing vehicles and their adjacent counterparts have been modeled using a Stackelberg game framework [10]. Such models enable adjustment of driving behaviors by varying emphasis on safety, efficiency, and comfort [15, 14]. However, Stackelberg games heavily rely on accurate predictions of surrounding vehicles driving styles, leading to significant errors when predictions are inaccurate. Additionally, Stackelberg games lack adaptability and flexibility in real driving, as real-world driving styles can deviate considerably from predefined styles.
@@
Figure 2: The causal evaluation-based evolutionary game theory for safe, efficienct, and cooperative interactive driving.
@@ IV Simulation Results
Figure 13: The reward curves of CEGT, Nash and Stackelberg across varied initial state settings over 10 runs for Case 1.
Figure 14: The speed curves for each AV using CEGT, Nash, and Stackelberg across varied initial state settings over 10 runs for Case 1.
@@ IV-A Simulation Case 1
Fig. 9 compares the average number of collisions for four algorithms: CEGT, EGT, Nash, and Stackelberg. CEGT demonstrates a significant advantage with an average collision count of approximately $1.5$, the lowest among all algorithms, compared to EGT ($2.8$), Nash ($4.5$), and Stackelberg ($3.6$). The data points for CEGT are tightly concentrated around the mean, reflecting consistent and reliable results, whereas Nash and Stackelberg show wider distributions with extreme outliers, reaching collision counts as high as $16$ and $18$, respectively. This stark contrast highlights the inefficiency of the alternative methods in ensuring safety. CEGT’s ability to minimize collisions and maintain stability across scenarios underscores its effectiveness in cooperative and adaptive decision-making.
@@
Fig. 10 illustrates the average TTC for CEGT, Nash, and Stackelberg over time, highlighting how well each algorithm balances safety and efficiency. CEGT maintains TTC values consistently above the safety threshold, ensuring a low risk of collisions throughout the simulation. In contrast, Nash and Stackelberg frequently fall below the safety threshold, particularly around $t=2~\text{s}$, where their TTC values drop as low as $300~\text{s}$, indicating a high risk of collisions. CEGT also avoids prolonged periods of negative TTC, which occur when the rear AV’s speed is slower than the lead AV’s, resulting in inefficiencies. Its TTC values remain mostly positive, demonstrating that it maintains efficient driving speeds. In comparison, Nash and Stackelberg often exhibit negative TTC values or positive values below the safety threshold for extended periods. For instance, at $t=5~\text{s}$, Nash has an average TTC value below $100~\text{s}$, indicating poor speed control. Additionally, CEGT demonstrates stable and consistent performance, with its TTC values generally staying above the safety threshold without significant fluctuations. In contrast, Nash and Stackelberg show far less stability, with frequent and severe drops in TTC, making their performance both unsafe and inefficient.
@@
Fig. 11 illustrates the average TTC between AV 1 and AV 2 for the CEGT, Nash, and Stackelberg algorithms. CEGT demonstrates superior performance by maintaining TTC values consistently close to or above the safety threshold for most of the simulation. CEGT maintains TTC values above safety threshold on average, ensuring low collision risk. In contrast, Nash and Stackelberg frequently exhibit TTC values below the safety threshold, with Nash dropping to as low as $80~\text{s}$, particularly around $t=2~\text{s}$, indicating high collision risks and inefficiency. CEGT also shows stable behavior, with fewer fluctuations compared to the other algorithms. This stability highlights its ability to manage interactions between AV 1 and AV 2 effectively. On the other hand, Nash and Stackelberg experience significant volatility, with frequent drops in TTC below $0~\text{s}$, reflecting poor speed control and increased risk of unsafe situations.
@@
Fig. 12 illustrates the average TTC between AV 3 and AV 4 for the CEGT, Nash, and Stackelberg algorithms. CEGT demonstrates superior performance by maintaining TTC values consistently above the safety threshold for most of the simulation, ensuring a low risk of collisions. In contrast, Nash and Stackelberg frequently fall below the safety threshold for extended periods, indicating higher collision risks. CEGT also avoids prolonged periods of negative TTC values, which occur when AV 3’s speed is slower than AV 4’s, leading to inefficiencies. Moreover, CEGT’s TTC values remain mostly positive, highlighting effective and efficient speed management. In comparison, Nash and Stackelberg exhibit significant volatility with frequent negative TTC values, reflecting poor coordination and inefficiencies.
@@
Fig. 13 The illustrates the total cumulative reward over time for the CEGT, Nash, and Stackelberg algorithms. CEGT clearly outperforms the other algorithms by achieving the highest cumulative reward throughout the simulation. At $t=10~\text{s}$, CEGT reaches a reward of approximately $162,388$, significantly higher than Nash around $49,545$ and Stackelberg about 176,039-176,039. This demonstrates CEGT’s superior ability to balance safety, efficiency, and cooperativeness. CEGT also shows a steeper growth rate in cumulative reward, reflecting its effective decision-making and efficient coordination among vehicles. In contrast, Nash has a much slower growth, while Stackelberg consistently exhibits a decline in cumulative reward, reflecting inefficient performance and frequent unsafe behaviors. Additionally, CEGT maintains a smaller variability range compared to Nash and Stackelberg, as shown by the narrower shaded region. This indicates more stable performance across different initial settings. In comparison, Nash and Stackelberg exhibit larger variability, highlighting their instability and less reliable outcomes.
@@
Fig. 14 illustrates the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stackelberg algorithms across 10 runs. CEGT demonstrates clear advantages in maintaining higher and more stable speeds compared to Nash and Stackelberg. For AV 1, CEGT has higher average speed than Nash and Stackelberg at most time points after $2~\text{s}$. For AV 2 to AV 4, CEGT achieves an average speed of approximately $10~\text{m/s}$ to $15~\text{m/s}$ throughout the simulation, indicating efficient driving behavior. In contrast, Nash and Stackelberg show significantly lower average speeds, with Stackelberg often around $5~\text{m/s}$ to $10~\text{m/s}$, and Nash below $5~\text{m/s}$, highlighting inefficiencies.
@@ IV-B Simulation Case 2
Fig. 17 demonstrates the average number of collisions for four algorithms: CEGT, EGT, Nash, and Stackelberg. CEGT demonstrates a significant advantage with an average collision count of approximately $0.4$, the lowest among all algorithms, compared to EGT around $2.5$, Nash around $3$, and Stackelberg around $0.7$. The data points for CEGT are tightly concentrated around the mean, reflecting consistent and reliable results, whereas Nash and Stackelberg show wider distributions with extreme outliers between $5$ to $10$. This stark contrast highlights the inefficiency of the alternative methods in ensuring safety. CEGT’s ability to minimize collisions and maintain stability across scenarios underscores its effectiveness in cooperative and adaptive decision-making.
@@
Fig. 18 illustrates the cumulative reward profiles over time for four AVs using the CEGT, Nash, and Stackelberg algorithms. For AV 1, CEGT shows a steady growth in cumulative reward, reaching approximately $0.7\times 10^{5}$ at $t=10~\text{s}$. This lower reward compared to other AVs is due to AV 1 prioritizing maintaining a similar distance to other AVs, thereby sacrificing its self-reward for the overall system’s benefit. For AV 2, the CEGT reward curve is similar to Stackelberg; however, CEGT exhibits narrower variation, suggesting greater stability and adaptability to different initial settings of AVs. Similar trends are observed for AVs 3 and 4, where CEGT achieves higher rewards than the other algorithms. Although Stackelberg also presents a consistent upward trend, its wide variability indicates poor capability to adapt to random initial settings of AVs and maintain reliable performance.
@@
Figure 18: The reward curves of CEGT, Nash and Stackelberg across varied initial state settings over 10 runs for Case 2.
@@
Figure 19: The speed curves for each AV using CEGT, Nash, and Stackelberg across varied initial state settings, with a collision penalty of 100-100, over 10 runs for Case 2.
@@
Figure 20: The speed curves for each AV using CEGT, Nash, and Stackelberg across varied initial state settings, with a collision penalty of 200-200, over 10 runs for Case 2.
@@
Figure 21: The speed curves for each AV using CEGT, Nash, and Stackelberg across varied initial state settings, with a collision penalty of 300-300, over 10 runs for Case 2.
@@
To verify the generalization of CEGT, different algorithm settings are also tested. Fig. 19 shows the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stackelberg algorithms across 10 runs, with a collision penalty of -100. CEGT demonstrates clear advantages in maintaining higher and more stable speeds compared to Nash and Stackelberg. For AV 1, CEGT achieves a lower average speed than Nash and Stackelberg at most time points due to its cooperation-oriented decision-making. This sacrifice by AV 1 allows other AVs to achieve significantly higher speeds. For AVs 2 to 4, CEGT maintains an average speed of approximately $10~\text{m/s}$ to $15~\text{m/s}$ throughout the simulation, reflecting efficient driving behavior. In contrast, Nash and Stackelberg show significantly lower average speeds, with Stackelberg remaining below $10~\text{m/s}$ and Nash often below $5~\text{m/s}$, highlighting inefficiencies.
@@
Fig. 20 shows the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stackelberg algorithms across 10 runs, with a collision penalty of -200. For AV 1, CEGT displays a lower average speed compared to Nash and Stackelberg at most time points, as it prioritizes cooperation over individual performance. This trade-off enables other AVs to achieve substantially higher speeds. For AVs 2 to 4, CEGT consistently maintains an average speed ranging from $10~\text{m/s}$ to $15~\text{m/s}$ throughout the simulation, showcasing its ability to ensure efficient driving behavior. In comparison, Nash and Stackelberg exhibit significantly reduced average speeds, with Stackelberg staying below $10~\text{m/s}$ and Nash frequently falling under $5~\text{m/s}$, underscoring their lack of efficiency.
@@
Fig. 21 shows the speed distributions with variations over time for four AVs using the CEGT, Nash, and Stackelberg algorithms across 10 runs, with a collision penalty of -300. The AV 1’s speed patterns show distinct characteristics across the three strategies. The CEGT approach maintains speeds between $8\text{-}12~\text{m/s}$ with moderate fluctuations, while demonstrating higher stability compared to other methods. The Nash equilibrium exhibits a sharp decline in speed after the initial period, settling at considerably low speeds around $2\text{-}3~\text{m/s}$. The Stackelberg approach achieves intermediate performance, maintaining speeds around $10\text{-}12~\text{m/s}$ with relatively consistent behavior. For AVs 2 through 4, similar patterns emerge but with notable differences. CEGT consistently demonstrates superior performance, maintaining speeds between $10\text{-}15~\text{m/s}$ across all vehicles, though with visible oscillations in the velocity profiles. The Nash equilibrium consistently performs poorly across all AVs, with speeds dropping dramatically to near-zero levels after the initial phase. The Stackelberg strategy shows intermediate performance but exhibits greater variance as indicated by the wider shaded regions. The shaded areas indicate that CEGT maintains relatively consistent behavior across multiple simulations, while Nash and Stackelberg strategies show higher variability. This suggests that CEGT not only achieves better average performance but also provides more reliable and predictable vehicle behavior.